Evian's average on five tests was 90%. The average of the four highest test scores was 93%, and the average of the four lowest test scores was 87.75%. What was the sum of the lowest mark and the highest mark?
Answer by Theo(13342)

your average mark for all 5 tests was 90%.

this was the sum of the 5 test scores divided by 5.

the total mark for all 5 tests was therefore 90 * 5 = 450%.

similarly the total mark for the 4 highest test scores was 4 * 93 = 372%

similarly, the total mark for the 4 lowest test scores was 4 * 87.75 = 351%

subtract the 4 lowest test scores from the total test score to get the highest test score.

subtract the 4 highest test scores from the total test score to get the lowest test score.

you get the highest test score = 450 - 351 = 99%

you get lowest test score = 450 - 372 = 78%

the total of the highest test score and the lowest test score is 177%.